OpenStudy (anonymous):

whats 5/8% as a decimal?

OpenStudy (anonymous):

5/(8/100) = (5 * 100)/8 = 500/8 = 62.5

OpenStudy (anonymous):

A. 0.625 B. 0.0625 C. 0.00625 D. 0.000625